OpenStudy (anonymous):

************** metal plzz help ************ use a special right triangle to write each trigonometric ratios as a fraction use this 30,60,90 45,45,90 1. cos60= 2. tan30= 3;.sin45=

OpenStudy (anonymous):

Use the unit circle.

OpenStudy (anonymous):

can you solve them

OpenStudy (anonymous):

1.= 1/2

OpenStudy (anonymous):

2.= \[\sqrt{3}\]

OpenStudy (anonymous):

3.= \[\sqrt{2}/2\]

OpenStudy (johnweldon1993):

Okay...a 30-60-90 triangle looks like  And a 45-45-90 triangle looks like